1. the index of refraction of a certain type of glass is 1.58 wht is the speed of of light in the glass ?
1
Expert's answer
2012-12-10T10:37:31-0500
The index of refraction, n, is a ratio that compares the speed of light in avacuum to the speed of light in another medium.
n = c/v where c = speed of light in a vacuum = 3 X 108 m/s

The speed of light in all other media is less than in a vacuum. Therefore, nis always larger than one. It may be thought of as the measure of the slowness of light in a medium: the slower light travels in medium, the larger n will be for that medium. Since n divides speed by speed it has no units

n = c/v
1.58 = (3 X 10^8 m/s)/v
v = 1.9 X 10^8 m/s